OpenStudy (anonymous):

Factor the polyomial completely using any method 3n^3-10n^2-48n+160 Medal given if answered

jhonyy9 (jhonyy9):

3n(n2-16)-10(n2-16)=(3n-10)(n2-16)

hero (hero):

n^2(3n-10)-16(3n-10) (3n-10)(n^2-16) (3n-10)(n+4)(n-4)

jhonyy9 (jhonyy9):

yes n2-16=(n-4)(n+4)

hero (hero):

jhonyy9, that's not completely

OpenStudy (anonymous):

thanks guys , medal gievn

OpenStudy (turingtest):

n=4 is a factor, so we will use synthetic division from here:

OpenStudy (turingtest):

I mean a zero

hero (hero):

Synthetic Division? For this? Not necessary

OpenStudy (turingtest):

(n-4)(3n^2+2n-40)...

OpenStudy (turingtest):

(n+4)(n-4)(3n-10) is the answer

OpenStudy (turingtest):

Man I'm lagging, I couldn't see any of your answers when I posted :/